what is the calculated answer for an 75.0 kg athlete running a "4-minute mile" (i.e. 4.00 min/mile)
in nm
1
Expert's answer
2020-09-30T04:17:39-0400

here given mass = 75 kg..... velocity = 4 min/mile = 0.25 miles/min = 0.25x1609.34/60 = 6.7m/s

wavelength =h/(mv) = (6.63*10^-34)/(75*6.7) = 1.32x10-36 =1.32x10-27 nm

so wavelength in nm = 1.32*10-27 nm Answer